# Geometric models of simple Lie algebras via singularity theory

### v2

@@ 1.4. Main result 1: Geometric root system
There is a categorification of this pairing. Namely, our pairing is the symmetrized Euler pairing of the Fukaya category of an isolated singularity being developed in [BCCJ23]. Also note that Fukaya–Seidel category gives a categorification of Seifert form somewhat indirectly. In [BCCJ23], only the morphism spaces of the category are defined, but their $A_{\infty}$-structures will be defined in the forthcoming work following the construction in [CCJ20].
@@
Let $M$ be the Milnor fiber of a simple curve singularity of type $\Gamma$. Then the following subsets of $H_{1}(M,\partial M;\mathbb{Z})$ are equal to each other.
@@
    $\{\alpha\in H_{1}(M,\partial M;\mathbb{Z})\;|\;$ $\alpha$ is a vanishing arc, namely $\mathrm{var}(\alpha)$ is a vanishing cycle }\}.
Given an embedded connected arc, one may ask when its geometric variation image becomes a vanishing cycle. We will find that (iii) provides a geometric criterion for this question. We remark that for any isolated curve singularity $f(x,y)$, geometric monodromy $\rho^{\prime}$ can be chosen so that $\rho^{\prime}$ does not have any fixed point away from the boundary.
The preimage of a vanishing cycle can be possibly a union of disjoint arcs, called vanishing arcsets in general. Such vanishing arcs and arcsets were introduced and studied in [BCC+25]. We refer readers to it for more details.
@@ 1.5. Main result 2: Coxeter wheels and geometric roots
The converse of the last statement holds for $A_{2k}$, $E_{7}$ and $E_{8}$ but not for others types. In fact the same phenomenon happens in Coxeter plane too. The projection of roots to the root projections in the Coxeter plane are one to one only for the above types and can have multiplicities in general. One can check that regarding oriented edges/spokes as vectors in $\mathbb{R}^{2}$ have exactly the same multiplicities.
@@
In the case of $A_{k}$ and $D_{k}$-wheels, they can be also related to Coxeter planes, but we will not do so because the wheel itself is simple enough. For the $A_{k}$-wheel given by the regular $(k+1)$-gon, number of oriented edges and diagonals between the vertices are exactly $2{k\choose 2}=k(k+1)$, they are exactly the geometric roots for $A_{k}$. For the $D_{k}$-wheel, opposite edges of 2(k1)2(k-1)-gon are identified, and from this, two parallel line segments that do not meet the center puncture and having the same orientation become homologous to each other. Number of such equivalence classes are ${2k-2\choose 2}-(k-1)$. Number of oriented spokes that meet the center is $2\times 2(k-1)$. In total, we have 2k(k1)2k(k-1) oriented edges/spokes, which equals the number of geometric roots.
@@ 1.6. Main result 3: Geometric construction of Lie algebra
    For $h\in H_{1}(M;\mathbb{Z}),\alpha\in\Phi_{\Gamma}$,
    $h,g_{\alpha}]=-[g_{\alpha},h]=(h\bullet\alpha+h\bullet\rho_{*}(\alpha))g_{\alpha}.$
    In particular,
@@
    For each geometric root $\alpha\in\Phi_{\Gamma}$,
    $g_{\alpha},g_{-\alpha}]=-\mathrm{var}(\alpha)\in H_{1}(M;\mathbb{Z}).$
@@
    For any two roots $\alpha,\beta\in\Phi_{\Gamma}$ such that $\alpha\neq-\beta$, we set
    $g_{\alpha},g_{\beta}]=\begin{cases}N_{\alpha,\beta}g_{\alpha+\beta}&\textrm{if }\;\alpha+\beta\in\Phi_{\Gamma},\\ 0&\textrm{if }\;\alpha+\beta\notin\Phi_{\Gamma}.\end{cases}$
    Here the sign $N_{\alpha,\beta}\in\{+1,-1\}$ is defined as
    $N_{\alpha,\beta}=(-1)^{\mathrm{var}(\beta)\bullet\alpha}.$
@@ 1.8. Main result 4: Description of Weyl group reflections and the Coxeter element
Theorem 1.22.
Within the Weyl group associated with $\Phi_{\Gamma}$, Coxeter element can be taken as $\overline{\rho}_{*}$, which is the orientation reversal composed with $\rho_{*}$.
@@ 5.6. Simple or Projective basis
We prove $\Phi_{\Gamma}^{(4)}\supseteq\Phi_{\Gamma}^{(3)}$ in two steps. First, we claim that given the projective basis $(\beta_{1},\ldots,\beta_{k})$, the collection of variation images of $(\beta_{k},\ldots,\beta_{1})$ forms a distinguished collection of vanishing cycles. Next, note that the monodromy image of a vanishing cycle is a vanishing cycle since one can compose the vanishing path with a loop going around all Morse critical points. In Section 5, we have shown that the monodromy action images of the projective basis cover all geometric roots, up to reversing orientation. Hence this proves $\Phi_{\Gamma}^{(4)}\supseteq\Phi_{\Gamma}^{(3)}$.
@@
For the simple basis $(\alpha_{1},\ldots,\alpha_{k})$, the matrix in the theorem is exactly $B_{ij}$ and we have checked that they are upper-triangular. This means that simple basis is in fact a distinguished collection of vanishing cycles. For the projective basis $(\beta_{1},\ldots,\beta_{k})$, one can similarly compute the matrix of variation operator and find that it is lower-triangular. Hence after reversing the ordering as $(\beta_{k},\ldots,\beta_{1})$, their variation images form a distinguished collection of vanishing cycles.
[…]
@@
Now it remains to prove Lemma 6.4 (3). For this, we will consider the set of vanishing cycles of $f(x,y)$ and compare it with that of its stabilization $\tilde{f}(x,y,z)=f(x,y)+z^{2}$.
@@
Let $f(x,y)$ be a simple curve singularity of type $\Gamma$, and $\Lambda_{f}\subset H_{1}(M;\mathbb{Z})$ be the set of vanishing cycles of $f$. Then we have $|\Lambda_{f}|\leq|\Phi_{\Gamma}|$. Since the variation operator is an isomorphism, it follows that $|\Phi_{\Gamma}^{(4)}|=|\Lambda_{f}|\leq|\Phi_{\Gamma}|$.
@@ 7. Construction of Lie algebras
    If $\alpha+\beta$ is non-zero and not a geometric root, then we get $[[A,B],C]=[[B,C],A]=[[C,A],B]=0$ and Jacobi identity holds. If $\alpha+\beta=0$, then
    $A,B],C]+[[B,C],A]+[[C,A],B]=0-(\gamma,-\alpha)(-\mathrm{var}(-\alpha))-(\gamma,\alpha)\mathrm{var}(\alpha)=0.$
    Assume that $\alpha+\beta\in\Phi_{\Gamma}$. Then Jacobi identity follows from the following identities:
@@ 8.1. $A_{2}$-case
The geometric Lie algebra from $A_{2}$-singularity $x^{3}+y^{2}$ is constructed as follows. The Milnor fiber $M$, defined by $x^{3}+y^{2}=1$, is a once-punctured torus (obtained by gluing the opposite edges of the punctured hexagon in Figure 24). The regular triangle formed by three punctures in Figure 24 is the $A_{2}$-wheel, and six geometric roots are homology classes of arcs $\alpha_{1},\alpha_{2},\alpha_{3}$ and those of their orientation reversals $\alpha_{1},-\alpha_{2},-\alpha_{3}$. Note that if we regard geometric roots as vectors, they are exactly the vectors in the $A_{2}$-root system.
Figure 24. Three roots $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$.
Monodromy images and variation images of geometric roots are illustrated in Figure 25. We take $\{\mathrm{var}(\alpha_{1}),\mathrm{var}(\alpha_{2})\}$ as the basis of $H_{1}(M;\mathbb{Z})$.
@@
Now the signs appearing in type (4) in the proof have the following geometric interpretation of the right hand rule. Lie bracket is non-zero if two roots can be added to another root. When it does, we need to determine the sign $\pm 1$. The claim is that for type $A$, the sign is given by the right hand rule. Namely, the Lie bracket $g_{\alpha_{1}},g_{\alpha_{2}}]=g_{\alpha_{3}}$ is realized by the vector sum $\alpha_{3}=\alpha_{1}+\alpha_{2}$ in Figure 24, and the sign is positive as the orientation given by the pair $(\alpha_{1},\alpha_{2})$ agrees with the (complex) orientation of the Milnor fiber. In contrast, the sign in the Lie braket $g_{-\alpha_{1}},g_{\alpha_{3}}]=-g_{\alpha_{2}}$ is negative as the pair $(-\alpha_{1},\alpha_{3})$ does not agree with the orientation. This sign rule holds only for $A_{k}$-cases.
@@ 8.3. $D_{k}$-case
$g_{\alpha_{1}},g_{\alpha_{2}}]=\begin{cases}-\epsilon_{\alpha_{1},\alpha_{2}}g_{\alpha_{1}+\alpha_{2}}&\textrm{if the triangle formed by}\;\alpha_{1},\alpha_{2},\mbox{ and }\alpha_{1}+\alpha_{2}\textrm{ contains}\;\;v_{0},\\ \epsilon_{\alpha_{1},\alpha_{2}}g_{\alpha_{1}+\alpha_{2}}&\textrm{otherwise.}\end{cases}$
@@
The sign is given by the parity of the intersection $\mathrm{var}(\alpha_{2})\bullet\alpha_{1}$. This can be checked through case by case analysis: first case is when the triangle $\alpha_{1},\alpha_{2},\alpha_{1}+\alpha_{2}$ does not meet $v_{0}$, second case is when the triangle contains $v_{0}$ as a vertex and last case is when the triangle contains $v_{0}$. We leave the details to the reader. ∎
Hence the sign is given by the right hand rule, with the additional (1)(-1) sign if three geometric roots form a triangle that contains the center $v_{0}$ in the interior. Signs for $E$-types are more subtle. One can specify the shape and size of the triangles formed by three vectors that requires additional (1)(-1) signs as in $D_{k}$-case, However, we omit it due to its intricacy.
@@ 9. Description of Weyl group reflections and a Coxeter element
In this section, we explain the Weyl group action on the set of geometric roots and find the relation between the Coxeter element and the monodromy of the singularity. Recall that our pairing is the negative symmetrized Seifert form. For a singularity of three variables, this is the same as intersection form, and $s_{\alpha}$ can be then identified with Picard–Lefschetz transformation, namely it comes from the Dehn twist with respect to the vanishing sphere $\mathrm{var}(\alpha)$. But for a singularity of two variables, the pairing is not the same as intersection form and $s_{\alpha}$ does not come from mapping class group of the Milnor fiber. We will see that it corresponds to certain “combinatorial operation” of flipping roots.
@@
Proposition 9.2.
The action of Coxeter element $c$ on geometric roots equals the operator $\overline{\rho}_{*}$: monodromy operator composed with an orientation reversal.
Remark 9.3.
Proposition 9.2 is a variant of a classical result concerning skew-symmetric Milnor lattices, as found on page 36 of [GZ95]. Nevertheless, we include a complete proof, as it plays a crucial role in the subsequent discussions.
Proof.
From Theorem 1.8, recall that the set $\Lambda_{f}$ of vanishing cycles forms a root system isomorphic to $\Phi_{\Gamma}$. Let $\sigma_{\Delta}$ denote the reflection on $H_{1}(M;\mathbb{Z})$ associated with the root $\Delta\in\Lambda_{f}$. The reflection operators on $H_{1}(M,\partial M;\mathbb{Z})$ are related with the reflection operators on $H_{1}(M;\mathbb{Z})$ by the formula
$s_{\alpha}=\mathrm{var}^{-1}\circ\sigma_{\mathrm{var}(\alpha)}\circ\mathrm{var}.$
From this, the Coxeter element of $\Phi_{\Gamma}$ can be expressed with a product of reflections in $\Lambda_{f}$:
$c=\mathrm{var}^{-1}\circ(\sigma_{\Delta_{1}}\circ\dots\circ\sigma_{\Delta_{k}})\circ\mathrm{var}.$
We will claim that $\sigma_{\Delta_{1}}\circ\dots\circ\sigma_{\Delta_{k}}$ is equal to the global monodromy operator on the absolute homology group $H_{1}(M;\mathbb{Z})$ composed with an orientation reversal. Then the proof of $c=\overline{\rho}_{*}$ on $H_{1}(M,\partial M;\mathbb{Z})$ follows, since $\rho_{*}$ and $\mathrm{var}$ commute.
Let $\tau_{\Delta}$ denote the Picard–Lefschetz transformation associated with the vanishing cycle $\tau_{\Delta}$. The composition $\tau_{\Delta_{1}}\circ\dots\circ\tau_{\Delta_{k}}$ of Picard–Lefschetz transforms is equal to the global monodromy $\rho_{*}$ on $H_{1}(M;\mathbb{Z})$, since $\{\Delta_{1},\dots,\Delta_{k}\}$ is a distinguished collection. Hence, we are going to show
(9.1) $S_{1}\dots S_{k}=-T_{1}\dots T_{k}$
where $S_{\lambda}$ is the matrix for the reflection $\sigma_{\Delta_{\lambda}}$ and $T_{\lambda}$ is the matrix for the Picard–Lefschetz transformation $\tau_{\Delta_{\lambda}}$ with respect to the basis $\{\Delta_{1},\dots,\Delta_{k}\}$ in $H_{1}(M;\mathbb{Z})$.
The matrix $S_{\lambda}$ is computed from the reflection formula
$\sigma_{\Delta_{\lambda}}(\Delta_{j})$ $=\Delta_{j}-(\Delta_{j},\Delta_{\lambda})\Delta_{\lambda}=\Delta_{j}-\big{(}B_{\lambda j}+B_{j\lambda}\big{)}\Delta_{\lambda}$
where $B$ is same with the matrix from Proposition 4.2. Hence, we get
$(S_{\lambda})_{ij}=\delta_{ij}-\delta_{i\lambda}\big{(}B_{\lambda j}+B_{j\lambda}\big{)}$
where $\delta_{ij}$ denotes the Kronecker delta. In other words,
$S_{\lambda}=\begin{pmatrix}1&&&&\\ &1&&&\\ &&\ddots&&\\ &&&1&\\ &&&&1\end{pmatrix}-\begin{pmatrix}0\\ \vdots\\ \text{$\lambda$-th row of $B+B^{t}$}\\ \vdots\\ 0\end{pmatrix}$.
The matrix $T_{\lambda}$ is computed from the (2-variable) Picard–Lefshetz formula
$\tau_{\Delta_{\lambda}}(\Delta_{j})$ $=\Delta_{j}-(\Delta_{j}\bullet\Delta_{\lambda})\Delta_{\lambda}=\Delta_{j}-\big{(}B_{\lambda j}-B_{j\lambda}\big{)}\Delta_{\lambda}.$
Here, the second equality follows from the relation between the Seifert form and the intersection form (see, e.g., [AGZV12] Theorem 2.4). Subsequently, we get
$T_{\lambda}=\begin{pmatrix}1&&&&\\ &1&&&\\ &&\ddots&&\\ &&&1&\\ &&&&1\end{pmatrix}-\begin{pmatrix}0\\ \vdots\\ \text{$\lambda$-th row of $B-B^{t}$}\\ \vdots\\ 0\end{pmatrix}$.
Note that $(S_{\lambda})_{\lambda,1}=-(T_{\lambda})_{\lambda,1},\dots,(S_{\lambda})_{\lambda,\lambda-1}=-(T_{\lambda})_{\lambda,\lambda-1}$ and $(S_{\lambda})_{\lambda,\lambda+1}=(T_{\lambda})_{\lambda,\lambda+1},\dots,(S_{\lambda})_{\lambda,k}=(T_{\lambda})_{\lambda,k}$ hold for every $\lambda$, as the matrix $B$ is upper-triangular. Moreover, $(S_{\lambda})_{\lambda,\lambda}=-(T_{\lambda})_{\lambda,\lambda}=-1$ holds for every $\lambda$, since all the diagonal entries of $B$ are $1$. Now the verification of (9.1) is a routine computation. The matrix identity below can be easily confirmed and extended to larger matrices.
$\begin{pmatrix}-a_{1}&a_{2}&a_{3}\\ 0&1&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&0\\ -b_{1}&-b_{2}&b_{3}\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&0\\ -c_{1}&-c_{2}&-c_{3}\end{pmatrix}=-\begin{pmatrix}a_{1}&a_{2}&a_{3}\\ 0&1&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&0\\ b_{1}&b_{2}&b_{3}\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&0\\ c_{1}&c_{2}&c_{3}\end{pmatrix}$.
In particular, Proposition 9.2, together with the discussions in Section 5, provides a geometric realization of Kostant’s theorem (Theorem 1.21) in terms of the Coxeter wheel.

### v3

@@ 1.4. Main result 1: Geometric root system
There is a categorification of this pairing. Namely, our pairing is the symmetrized Euler pairing of the Fukaya category $\mathcal{F}_{\rho}$ of an isolated singularity being developed in [BCCJ23]. Also note that Fukaya–Seidel category gives a categorification of Seifert form somewhat indirectly. In [BCCJ23], only the morphism spaces of the category are defined, but their $A_{\infty}$-structures will be defined in the forthcoming work following the construction in [CCJ20].
@@
Let $M$ be the Milnor fiber of a simple curve singularity of type $\Gamma$. Then the following subsets of $H_{1}(M,\partial M;\mathbb{Z})$, which are denoted by $\Phi^{(1)}_{\Gamma},\ldots,\Phi^{(4)}_{\Gamma}$, are equal to each other.
@@
    $\{\alpha\in H_{1}(M,\partial M;\mathbb{Z})\;|\;$ $\mathrm{var}(\alpha)$ is a member of a distinguished collection of vanishing cycles of $f$ }\}.
(iii) and (iv) provide two geometric characterization of roots. First, (iii) provides an easy geometric way to determine whether $\alpha$ is a root or not using the monodromy $\rho$. We remark that for any isolated curve singularity $f(x,y)$, geometric monodromy $\rho^{\prime}$ can be chosen so that $\rho^{\prime}$ does not have any fixed point away from the boundary.
(iv) demonstrates how the conventional notion that vanishing cycles are roots is also connected to the current approach. But let us point out that we are not taking all vanishing cycles, but rather the proper subset given by the members of distinguished collection of vanishing cycles.
This difference is largely because we are working with 2-variable rather 3-variable singularities (c.f. [HL24]). Roughly speaking, a vanishing cycle from an embedded vanishing path belongs to $\Phi^{(4)}_{\Gamma}$. But general vanishing cycles may come from immersed vanishing paths (due to the Picard–Lefschetz transformation). To distinguish them, we call the former em vanishing cycle.
In [GZ80], it is proven that every vanishing cycle of a 3-variable simple singularity is an em vanishing cycle, but this fails to hold for 2-variables. In this case the set of all vanishing cycles are much bigger than that of em vanishing cycles.
The geometric condition $L\cap\rho(L)=\emptyset$ (iii) has its origin in [BCC+25], where the notion of vanishing arcs and arcsets are introduced. Under the assumption that it is not A, D type, and $g>4$, this condition was sufficient to guarantee the arc to be a geometric vanishing arc (not necessarily em type). Since this assumption is false in our cases, such a finding is not applicable here.
@@ 1.5. Main result 2: Coxeter wheels and geometric roots
For the case of $A_{2}$ root system, $A_{2}$-Milnor fiber is a regular hexagon with 3 punctures (with opposite sides identified) as in Figure 2. $A_{2}$-wheel is a regular triangle whose vertices are these punctures. There are six geometric roots given by $\alpha_{1},\alpha_{2},\alpha_{3}$ and their orientation reversals (whose homology classes are $\alpha_{1},-\alpha_{2},-\alpha_{3}$).
(a) Classical roots.
(b) Geometric roots.
Figure 2. $A_{2}$ root systems from Lie theory and geometry.
In this case, the classical $A_{2}$ root system lies in $\mathbb{R}^{2}$, and the identification is almost immediate.
@@
The converse of the last statement of the theorem holds for $A_{2k}$, $E_{7}$ and $E_{8}$ but not for others types. In fact the same phenomenon happens in Coxeter plane too. The projection of roots to the root projections in the Coxeter plane are one to one only for the above types and can have multiplicities in general. One can check that regarding oriented edges/spokes as vectors in $\mathbb{R}^{2}$ have exactly the same multiplicities.
@@
In the case of $A_{k}$ and $D_{k}$-wheels, they can be also related to Coxeter planes, but we will not do so because the wheel itself is simple enough. For the $A_{k}$-wheel given by the regular $(k+1)$-gon, number of oriented edges and diagonals between the vertices are exactly $2{k\choose 2}=k(k+1)$, they are exactly the geometric roots for $A_{k}$.
For the $D_{k}$-wheel, opposite edges of 2(k1)2(k-1)-gon are identified, and from this, two parallel line segments that do not meet the center puncture and having the same orientation become homologous to each other. Number of such equivalence classes are ${2k-2\choose 2}-(k-1)$. Number of oriented spokes that meet the center is $2\times 2(k-1)$. In total, we have 2k(k1)2k(k-1) oriented edges/spokes, which equals the number of geometric roots.
@@ 1.6. Main result 3: Geometric construction of Lie algebra
    For $h\in H_{1}(M;\mathbb{Z}),\alpha\in\Phi_{\Gamma}$, [h, g_α] = -[g_α,h] = ( h α+ h ρ_*(α)) g_α. In particular,
@@
    For each geometric root $\alpha\in\Phi_{\Gamma}$, [g_α, g_-α] = - var(α) H_1(M;Z).
@@
    For any two roots $\alpha,\beta\in\Phi_{\Gamma}$ such that $\alpha\neq-\beta$, we set [g_α, g_β] = {Nα,βgα+βif   α+ βΦΓ,0 if  α+ βΦΓ. Here the sign $N_{\alpha,\beta}\in\{+1,-1\}$ is defined as N_α,β = (-1)^var(β) α.
@@
Remark 1.18.
There are well-known categorical constructions of Lie agebras by Peng-Xiao [PX97], [PX00] via Hall algebras based on the works of Ringel [Rin90a], [Rin90b]. It would be very interesting to find a connection to our work. As an evidence, in the case of $A$, all Lie brackets of type $g_{\alpha},g_{\beta}]=g_{\gamma}$, the arc for $\gamma$ is the unique extension of arc for $\alpha$ by arc for $\beta$ in the Fukaya category $\mathcal{F}_{\rho}$.
@@ 1.8. Main result 4: Description of Weyl group reflections and the Coxeter element
Recall from Proposition 2.4 of Hubery–Krause [HK16] that Coxeter element equals the negative of the monodromy. This equals $\overline{\rho}_{*}$, which is the orientation reversal composed with $\rho_{*}$.
@@ 5.6. Simple or Projective basis
We prove $\Phi_{\Gamma}^{(4)}\supseteq\Phi_{\Gamma}^{(3)}$ in two steps. First, we claim that given the projective basis $(\beta_{1},\ldots,\beta_{k})$, the collection of variation images of $(\beta_{k},\ldots,\beta_{1})$ forms a distinguished collection of vanishing cycles. Next, note that the monodromy image of an em vanishing cycle is an em vanishing cycle since one can compose the vanishing path with a loop going around all Morse critical points. In Section 5, we have shown that the monodromy action images of the projective basis cover all geometric roots, up to reversing orientation. Hence this proves $\Phi_{\Gamma}^{(4)}\supseteq\Phi_{\Gamma}^{(3)}$.
@@
For the simple basis $(\alpha_{1},\ldots,\alpha_{k})$, the matrix in the theorem is exactly $B_{ij}$ and we have checked that they are upper-triangular. This means that simple basis is in fact a distinguished collection of vanishing cycles. For the projective basis $(\beta_{1},\ldots,\beta_{k})$, one can similarly compute the matrix of variation operator and find that it is lower-triangular. Hence after reversing the ordering as $(\beta_{k},\ldots,\beta_{1})$, their variation images form a distinguished collection of vanishing cycles. ∎
@@
Now it remains to prove Lemma 6.4 (3). For this, we will consider the set of em vanishing cycles of $f(x,y)$ and compare it with that of its stabilization $\tilde{f}(x,y,z)=f(x,y)+z^{2}$.
@@
Let $f(x,y)$ be a simple curve singularity of type $\Gamma$, and $\Lambda_{f}\subset H_{1}(M;\mathbb{Z})$ be the set of em vanishing cycles of $f$. Then we have $|\Lambda_{f}|\leq|\Phi_{\Gamma}|$. Since the variation operator is an isomorphism, it follows that $|\Phi_{\Gamma}^{(4)}|=|\Lambda_{f}|\leq|\Phi_{\Gamma}|$.
@@ 7. Construction of Lie algebras
    If $\alpha+\beta$ is non-zero and not a geometric root, then we get $[[A,B],C]=[[B,C],A]=[[C,A],B]=0$ and Jacobi identity holds. If $\alpha+\beta=0$, then [[A,B],C]+ [[B,C],A]+[[C,A],B]= 0 -(γ,-α) (-var(-α)) - (γ,α) var(α) =0. Assume that $\alpha+\beta\in\Phi_{\Gamma}$. Then Jacobi identity follows from the following identities:
@@ 8.1. $A_{2}$-case
The geometric Lie algebra from $A_{2}$-singularity $x^{3}+y^{2}$ is constructed as follows. The Milnor fiber $M$, defined by $x^{3}+y^{2}=1$, is a once-punctured torus (obtained by gluing the opposite edges of the punctured hexagon in Figure 2) (B). The regular triangle formed by three punctures in Figure 2 (B) is the $A_{2}$-wheel, and six geometric roots are homology classes of arcs $\alpha_{1},\alpha_{2},\alpha_{3}$ and those of their orientation reversals $\alpha_{1},-\alpha_{2},-\alpha_{3}$. Monodromy images and variation images of geometric roots are illustrated in Figure 25. We take $\{\mathrm{var}(\alpha_{1}),\mathrm{var}(\alpha_{2})\}$ as the basis of $H_{1}(M;\mathbb{Z})$.
@@
Now the signs appearing in type (4) in the proof have the following geometric interpretation of the right hand rule. Lie bracket is non-zero if two roots can be added to another root. When it does, we need to determine the sign $\pm 1$. The claim is that for type $A$, the sign is given by the right hand rule. Namely, the Lie bracket $g_{\alpha_{1}},g_{\alpha_{2}}]=g_{\alpha_{3}}$ is realized by the vector sum $\alpha_{3}=\alpha_{1}+\alpha_{2}$ in Figure 2 (B), and the sign is positive as the orientation given by the pair $(\alpha_{1},\alpha_{2})$ agrees with the (complex) orientation of the Milnor fiber. In contrast, the sign in the Lie braket $g_{-\alpha_{1}},g_{\alpha_{3}}]=-g_{\alpha_{2}}$ is negative as the pair $(-\alpha_{1},\alpha_{3})$ does not agree with the orientation. This sign rule holds only for $A_{k}$-cases.
@@ 8.3. $D_{k}$-case
(a) $g_{\alpha_{1}},g_{\alpha_{2}}]=-g_{\alpha_{1}+\alpha_{2}}$ since $\epsilon=-1$.
(b) $g_{\alpha_{1}},g_{\alpha_{2}}]=-(-1)g_{\alpha_{1}+\alpha_{2}}$.
Figure 26. Illustration of $D_{k}$ Lie bracket in Proposition 8.4.
@@
[…]
@@
The sign is given by the parity of the intersection $\mathrm{var}(\alpha_{2})\bullet\alpha_{1}$. This can be checked through case by case analysis: first case is when the triangle $\alpha_{1},\alpha_{2},\alpha_{1}+\alpha_{2}$ does not meet $v_{0}$, second case is when the triangle contains $v_{0}$ as a vertex, and last case is when the triangle contains $v_{0}$. We leave the details to the reader. ∎
Hence the sign is given by the right hand rule, with the additional (1)(-1) sign if three geometric roots form a triangle that contains the center $v_{0}$ in the interior. Signs for $E$-types are more subtle. One can specify the shape and size of the triangles formed by three vectors that requires additional (1)(-1) signs as in $D_{k}$-case, but omit them as there are too many to list.
@@ 9. Description of Weyl group reflections and a Coxeter element
In this section, we explain the Weyl group action on the set of geometric roots and find the relation between the Coxeter element and the monodromy of the singularity. Recall that our pairing is the negative symmetrized Seifert form. For a simple singularity of three variables, this is the same as intersection form, and $s_{\alpha}$ can be then identified with Picard–Lefschetz transformation. In particular, the corresponding Weyl group can be interpreted as the monodromy group. For a simple singularity of two variables, on the other hand, the pairing is not the same as intersection form and $s_{\alpha}$ cannot be then identified with Picard–Lefschetz transformation, so we need an alternative description for the Weyl group. We will see that it corresponds to certain “combinatorial operation” of flipping roots.
@@
Proposition 2.4 of Hubery–Krause [HK16] says that the action of Coxeter element $c$ on geometric roots equals the operator $\overline{\rho}_{*}$, monodromy operator composed with an orientation reversal. This is a variant of a classical result concerning skew-symmetric Milnor lattices, as found on page 36 of [GZ95].
In particular, this together with the discussions in Section 5, provides a geometric realization of Kostant’s theorem (Theorem 1.22) in terms of the Coxeter wheel.